# Semantic Communication-Enabled Wireless Adaptive Panoramic Video Transmission

### v1

@@ II-B WA Module
To measure the quality of panoramic video in the observation space more accurately, the evaluation metrics WS-PSNR and WS-SSIM take into account the panoramic frames structure information [13, 14]. These metrics combine the distortion relationship between projection planes and user’s field of view to map the spherical distortion to the 2D plane distortion, which can describe the immersive subjective experience quality more accurately in viewing panoramic video.
@@
$\mathbf{M}_{t}$ $\!=\!\sigma\!\left(\mathrm{Conv}_{3\times 3}^{2}\left\{\mathrm{AvgPool}^{4}\left(\mathbf{w}\!\right);\mathrm{MaxPool}^{4}\left(\mathbf{w}\right);\mathbf{y}_{t}\right\}\right)$
$\!=\!\sigma\left(\mathrm{Conv}_{3\times 3}^{2}\left\{\mathbf{F}_{\mathrm{Avg}};\mathbf{F}_{\mathrm{Max}};\mathbf{y}_{t}\right\}\right),$ (3)
where $\mathrm{AvgPool}^{4}$ and $\mathrm{MaxPool}^{4}$ denote that the initial weight maps $\mathbf{w}=\{w_{ij}\}\in\mathbb{R}^{H\times W}$ is downsampled by successive average pooling and maximum pooling four times. $H$ and $W$ represent the height and width of panoramic frames, respectively. Besides, a dimension is inserted into the pooling results to represent the number of channels, so the generated feature weight distributions are $\mathbf{F}_{\mathrm{Avg}}\in\mathbb{R}^{1\times\frac{H}{16}\times\frac{W}{16}}$ and $\mathbf{F}_{\mathrm{Max}}\in\mathbb{R}^{1\times\frac{H}{16}\times\frac{W}{16}}$, respectively. $\mathrm{Conv}_{3\times 3}^{2}$ denotes two successive convolution operations with convolution kernel of $3\times 3$, and $\sigma$ is the Sigmoid activation function. Besides, the weight values decrease gradually from the equator to the poles and the distribution of $w_{ij}$ is given as
$w_{ij}=\cos{\left(\left(i-\frac{H}{2}+\frac{1}{2}\right)\times\frac{\pi}{H}\right)}.$ (4)
Similar to the primary link, the weighted spatial attention map $\mathbf{M}_{t}^{\mathrm{ml}}$ in motion link can be obtained as
$\mathbf{M}_{t}^{\mathrm{ml}}$ $\!=\!\sigma\left(\mathrm{Conv}^{3\times 3}\left\{\mathrm{AvgPool}^{4}\left(\mathbf{w}\right)\!;\!\mathrm{MaxPool}^{4}\left(\mathbf{w}\right)\!;\!\mathbf{y}_{t}^{\mathrm{ml}}\right\}\right)$ (5)
$\!=\!\sigma\left(\mathrm{Conv}^{3\times 3}\left\{\mathbf{F}_{\mathrm{Avg}};\mathbf{F}_{\mathrm{Max}};\mathbf{y}_{t}^{\mathrm{ml}}\right\}\right).$
The feature information after deflation is obtained by multiplying the feature map with the points corresponding to the weighted spatial attention map. This process can be described as ${\mathbf{y^{\prime}}}_{t}^{\mathrm{ml}}=\mathbf{y}_{t}^{\mathrm{ml}}\otimes\mathbf{M}_{t}^{\mathrm{ml}}$ in the motion link and $\mathbf{y}_{t}^{\prime}=\mathbf{y}_{t}\otimes\mathbf{M}_{t}$ in the primary link, respectively.
@@
Figure 2: Network architectures of the primary link. $k\times k$ Conv is a convolution with $k\times k$ filters, and the output channels of convolution are given on horizontal line. $\uparrow 2$ and $\downarrow 2$ indicate upsampling and downsampling with a stride of 2. GDN denotes the Generalised Divisive Normalization in [19], IGDN denotes the inverse operation of GDN.
II-C Latitude Adaptive Module
@@
To reduce the information redundancy of transmission and improve the transmission efficiency of panoramic video, we propose the latitude adaptive network. For such a network, we design the adaptive weight feature map $\boldsymbol{\omega}_{t}=\left\{\omega_{t,mn}\right\}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}}$, which can be expressed as
$\omega_{t,mn}=\eta_{t,mn}{w}^{\mathrm{AAP}}_{mn}+1-\eta_{t,mn},$ (6)
where $\omega_{t,mn}$ denotes the adaptive weight feature value corresponding to the feature point $\left(m,n\right)$ at moment $t$. ${w}^{\mathrm{AAP}}_{mn}$ denotes the weight value of $\mathbf{w}$ after the adaptive average pooling (AAP) operation. $\eta_{t,mn}$ denotes the adaptive weight factor value of weight factor map $\boldsymbol{\eta}_{t}=\left\{\eta_{t,mn}\right\}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}}$, which is learned by the network according to the magnitude of entropy information. It can be expressed as $\boldsymbol{\eta}_{t}=f_{\mathrm{AW}}\left(\mathbf{e}_{t}\right)$, where $f_{\mathrm{AW}}$ denotes the function of adaptive weight factor network and $\mathbf{e}_{t}=\left\{e_{t,mn}\right\}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}}$ means the entropy map corresponding to the feature map.
In order to achieve adaption of feature information dimension in latitude, the dimensions of the information in the wireless channel cannot exceed the limit of the maximum dimension given in the corresponding latitude. This constraint can be expressed as
$l_{t,mn}\leq\omega_{t,mn}\times\mathrm{max}\left(\mathcal{Q}\right),$ (7)
where $l_{t,mn}$ denotes the information dimension corresponding to feature point $\left(m,n\right)$ by quantizing, which can be expressed as the number of channels. $\mathrm{max}\left(\mathcal{Q}\right)$ is the maximum value of the quantized set, which is given in Section IV. Therefore, the loss function of this module is expressed as
$\mathcal{L}_{t}^{\mathrm{la}}=\sum_{m}\sum_{n}{\mathrm{max}\left(0,\omega_{t,mn}\times\mathrm{max}\left(\mathcal{Q}\right)-l_{t,mn}\right)}.$ (8)
@@
=$\mathcal{L}^{\prime}{{}_{t}^{\mathrm{la}}}=\sum_{m}\sum_{n}{\mathrm{max}\left(0,\omega_{t,mn}\times\mathrm{max}\left(\mathbf{e}_{t}\right)-e_{t,mn}\right)}.$ (9)
@@
II-D Optimization Goal
@@
The APVST network aims to maximize the transmission quality of frames with the minimum channel bandwidth overhead. In fact, the frames are transmitted continuously, $N$ consecutive frames are one group of pictures (GoP). In the training phase, as shown in Fig. 1, the previously recovered panoramic frame is used as the reference frame for the current frame to achieve panoramic video transmission. Therefore, the loss function is formulated as follows:
@@
$\mathcal{L}\!=\!\frac{1}{N}\sum_{t=1}^{N}\left({D_{t}\!+\!\alpha\sum_{m}\sum_{n}{\left(e_{t,mn}\!+\!e_{t,mn}^{\mathrm{ml}}\right)}\!+\!\beta\mathcal{L}^{\prime}{{}_{t}^{\mathrm{la}}}}\right),$ (10)
@@
where $D_{t}$ represents the magnitude of image distortion, which is chosen as the weighted mean squared error (WMSE) for WS-PSNR, and as the inverse of WS-SSIM for WS-SSIM to indicate the distortion of frames [13, 14]. To minimize channel bandwidth overhead, we take the entropy $e_{t,mn}$ in the primary link and $e_{t,mn}^{\mathrm{ml}}$ in motion link into account, which can be expressed as the channel bandwidth cost. $\alpha$ and $\beta$ denote the balance coefficients, which are constants.
III Internal Structure of APVST
@@
In this section, we introduce the internal structure of the APVST network. The disassembled parts of the internal structure are shown in Fig. 2 and Fig. 3. Since the structures of primary and motion links are similar, the primary link will be introduced in detail.
III-A The Overall Structure of APVST
As shown in Fig. 2, the overall structure of APVST includes semantic recovery and extraction as well as Deep JSCC encoding and decoding.

### v2

@@ II-B1 WA module
To measure the quality of panoramic video in the observation space more accurately, the evaluation metrics WS-PSNR and WS-SSIM take into account the panoramic frames structure information [22], [23]. These metrics combine the distortion relationship between projection planes and the user’s FoV to map the spherical distortion to the 2D plane distortion, which can describe the immersive subjective experience quality more accurately in viewing the panoramic video. The specific definition can be expressed as
WS-PSNR $=10\log{\left(\frac{\text{MAX}}{\text{WMSE}}\right)}(\text{dB}),$
WS-SSIM $=-10\log{\left(1-\frac{\sum_{k=0}^{K}\text{WSSIM}_{k}}{\sum_{k=0}^{K}{\rm{w}}_{k}^{c}}\right)}(\text{dB}),$
where MAX represents the maximum value among all pixel points. $K$ means that the frame is divided into $K$ windows and ${\text{w}}_{k}^{c}$ is the weight of the center point of $k$-th sliding window. WMSE and $\text{WSSIM}_{k}$ can be expressed as
WMSE $=\frac{\sum_{i=1}^{H}\sum_{j=1}^{W}\left[\left({\hat{x}}(i,j)-x(i,j)\right)^{2}\times w(i,j)\right]}{\sum_{i=1}^{H}\sum_{j=1}^{W}w(i,j)},$
$\text{WSSIM}_{k}$ $=\left(\frac{2\mu_{\rm{x}_{k}}\mu_{\hat{\rm{x}}_{k}}+c_{1}}{\mu_{\rm{x}_{k}}^{2}+\mu_{\hat{\rm{x}}_{k}}^{2}+c_{1}}\right)\left(\frac{2\sigma_{{\rm{x}_{k}}\hat{\rm{x}}_{k}}+c_{2}}{\sigma_{\rm{x}_{k}}^{2}+\sigma_{\hat{\rm{x}}_{k}}^{2}+c_{2}}\right)\times{\rm{w}}_{k}^{c},$
where the frame size is $H\times W$. ${\hat{x}}(i,j)$ and $x(i,j)$ represent the pixel values at coordinates $\left(i,j\right)$ in the reconstructed and the original frames, respectively. $\mu$ and $\sigma$ represent the mean and variance, respectively. $c_{1}$ and $c_{2}$ are constants. Besides, the weight values $w_{ij}$ decrease gradually from the equator to the poles, and the distribution of $w_{ij}$ is given as
$w(i,j)=\cos{\left(\left(i-\frac{H}{2}+\frac{1}{2}\right)\times\frac{\pi}{H}\right)}.$ (17)
@@
$\mathbf{M}_{u,t}$ $\!=\!\sigma\!\left(\text{Conv}_{3\times 3}^{2}\left\{\text{AvgPool}^{4}\left(\mathbf{w}\!\right);\text{MaxPool}^{4}\left(\mathbf{w}\right);\mathbf{y}_{u,t}\right\}\right)$
$\!=\!\sigma\left(\text{Conv}_{3\times 3}^{2}\left\{\mathbf{F}_{\text{Avg}};\mathbf{F}_{\text{Max}};\mathbf{y}_{u,t}\right\}\right),$ (18)
where $\text{AvgPool}^{4}$ and $\text{MaxPool}^{4}$ denote that the initial weight maps $\mathbf{w}=\{w(i,j)\}\in\mathbb{R}^{H\times W}$ is downsampled by successive average pooling and maximum pooling four times. $H$ and $W$ represent the height and width of panoramic frames, respectively. Besides, a dimension is inserted into the pooling results to represent the number of channels, so the generated feature weight distributions are $\mathbf{F}_{\text{Avg}}\in\mathbb{R}^{1\times\frac{H}{16}\times\frac{W}{16}}$ and $\mathbf{F}_{\text{Max}}\in\mathbb{R}^{1\times\frac{H}{16}\times\frac{W}{16}}$, respectively. $\text{Conv}_{3\times 3}^{2}$ denotes two successive convolution operations with convolution kernel of $3\times 3$, and $\sigma$ is the Sigmoid activation function.
Similar to the primary link, the weighted spatial attention map $\mathbf{M}_{t}^{\text{ml}}$ in motion link can be obtained as
$\mathbf{M}_{u,t}^{\text{ml}}$ $\!=\!\sigma\left(\text{Conv}^{3\times 3}\left\{\text{AvgPool}^{4}\left(\mathbf{w}\right)\!;\!\text{MaxPool}^{4}\left(\mathbf{w}\right)\!;\!\mathbf{y}_{u,t}^{\text{ml}}\right\}\right)$ (19)
$\!=\!\sigma\left(\text{Conv}^{3\times 3}\left\{\mathbf{F}_{\text{Avg}};\mathbf{F}_{\text{Max}};\mathbf{y}_{u,t}^{\text{ml}}\right\}\right).$
The feature information after deflation is obtained by multiplying the feature map with the points corresponding to the weighted spatial attention map. This process can be described as ${\mathbf{y^{\prime}}}_{u,t}^{\text{ml}}=\mathbf{y}_{u,t}^{\text{ml}}\otimes\mathbf{M}_{u,t}^{\text{ml}}$ in the motion link and $\mathbf{y}_{u,t}^{\prime}=\mathbf{y}_{u,t}\otimes\mathbf{M}_{u,t}$ in the primary link, respectively.
@@
II-B2 Latitude Adaptive Module
@@
To reduce the information redundancy of transmission and improve the transmission efficiency of panoramic video, we propose the latitude adaptive network. For such a network, we design the adaptive weight feature map $\boldsymbol{\omega}_{u,t}=\left\{\omega_{u,t}(m,n)\right\}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}}$, which can be expressed as
$\omega_{u,t}(m,n)=\eta_{u,t}(m,n){w}_{\text{AAP}}(m,n)+1-\eta_{u,t}(m,n),$ (20)
where $\omega_{u,t}(m,n)$ denotes the adaptive weight feature value corresponding to the feature point $\left(m,n\right)$ at time $t$. ${w}_{\text{AAP}}(m,n)$ denotes the weight value of $\mathbf{w}$ after the adaptive average pooling (AAP) operation. $\eta_{u,t}(m,n)$ denotes the adaptive weight factor value of weight factor map $\boldsymbol{\eta}_{u,t}=\left\{\eta_{u,t}(m,n)\right\}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}}$, which is learned by the network according to the magnitude of entropy information. It can be expressed as $\boldsymbol{\eta}_{u,t}=f_{\text{AW}}\left(\mathbf{e}_{u,t}\right)$, where $f_{\text{AW}}$ denotes the function of adaptive weight factor network and $\mathbf{e}_{u,t}=\left\{e_{u,t}(m,n)\right\}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}}$ means the entropy map corresponding to the feature map.
In order to achieve the adaption of feature information dimension in latitude, the dimensions of the information in the wireless channel cannot exceed the limit of the maximum dimension given in the corresponding latitude. This constraint can be expressed as
$l_{u,t}(m,n)\leq\omega_{u,t}(m,n)\times{\max}\left(\mathcal{Q}\right),$ (21)
where $l_{u,t}(m,n)$ denotes the information dimension corresponding to feature point $\left(m,n\right)$ by quantizing, which can be expressed as the number of channels. ${\max}\left(\mathcal{Q}\right)$ is the maximum value of the quantized set, which is given in Section V. Therefore, the loss function of this module is expressed as
$\mathcal{L}_{u,t}^{\text{la}}=\sum_{m}\sum_{n}{{\max}\left(0,\omega_{u,t}(m,n)\times{\max}\left(\mathcal{Q}\right)-l_{u,t}(m,n)\right)}.$ (22)
@@
=$\mathcal{L}^{\prime}{{}_{u,t}^{\text{la}}}\!=\!\sum_{m}\sum_{n}{{\max}\left(0,\omega_{u,t}(m,n)\!\times\!{\max}\left(\mathbf{e}_{u,t}\right)\!-\!e_{u,t}(m,n)\right)}.$ (23)
@@
II-C Problem Formulation
For the above description of the system model, we decouple it into two problems and optimize them using different objective functions and methods. The optimization goal for the semantic transmission network APVST is to maximize users’ immersive experience quality with minimal bandwidth consumption. Meanwhile, the optimization goal for RSMA is to maximize all users’ QoS scores based on semantic communication.
II-C1 The optimization goal of APVST
@@
The APVST network aims to maximize the transmission quality of frames with the minimum channel bandwidth overhead. In fact, the frames are transmitted continuously, $N$ consecutive frames are one group of pictures (GoP). In the training phase, the previously recovered panoramic frame is used as the reference frame for the current frame to achieve panoramic video transmission. Therefore, the loss function is formulated as
@@
$\mathcal{L}_{u}\!=\!\frac{1}{N}\!\sum_{t=1}^{N}\left[{D\left(\mathbf{x}_{u,t},\hat{\mathbf{x}}_{u,t}\right)\!+\!\alpha\!\left(\mathcal{L}_{u,t}^{\text{enml}}\!+\!\mathcal{L}_{u,t}^{\text{en}}\!\right)\!+\!\beta\mathcal{L}^{\prime}{{}_{u,t}^{\text{la}}}}\right],$ (24)
@@
where $D\left(\mathbf{x}_{u,t},\hat{\mathbf{x}}_{u,t}\right)$ represents the magnitude of image distortion, which is chosen as the WMSE for WS-PSNR, and as the inverse of WS-SSIM for WS-SSIM to indicate the distortion of frames [22], [23]. To minimize channel bandwidth overhead, we take the entropy $e_{u,t}(m,n)$ and $e_{u,t}^{\text{ml}}(m,n)$ into account. $\mathcal{L}_{u,t}^{\text{enml}}=\sum_{m}\sum_{n}{e_{u,t}^{\text{ml}}(m,n)}$ and $\mathcal{L}_{u,t}^{\text{en}}=\sum_{m}\sum_{n}{e_{u,t}(m,n)}$ represent the channel bandwidth cost of motion link and primary link, respectively. $\alpha$ and $\beta$ denote the balance coefficients, which are constants. In section III, the internal structure of the APVST network will be introduced, and the optimization of the network is completed based on this loss function.
II-C2 Problem Formulation of RSMA-enabled semantic information transmission
In the process of panoramic video transmission, the factors that most significantly affect the QoS are the transmission delay and the quality of the immersive experience for the users. To comprehensively consider these two components, QoS is transformed into a scoring model which is decomposed into two parts: the transmission delay score and the immersive experience quality score.
To better represent the data size of the semantic codeword sequence to be transmitted at time slot $t$, referring to [26], let $\lambda$ denote the number of bytes per semantic codeword. The data size for user $u$ at time slot $t$ is given as
$D_{u,t}=\lambda k_{u,t},$ (25)
where $k_{u,t}$ represents the dimension of the semantic codeword sequence $\mathbf{s}_{u,t}$ for user $u$ at time slot $t$, which can be adjusted by selecting different values from the quantization set $\mathcal{Q}$. Generally speaking, the quality of the user’s immersive experience increases with the increase of channel bandwidth ratio (CBR) [27], which can be expressed as $\frac{k_{u,t}}{H\times W\times 3}$, where 3 represents the original panoramic frame has three RGB channels. The transmission delay of user $u$ at time slot $t$ can be expressed as
$T_{u,t}=\frac{D_{u,t}}{R_{u,t}}=\frac{\lambda k_{u,t}}{R_{u,t}^{p}+C_{u,t}}.$ (26)
Transmission delay plays a key role in users’ QoS, yet it exhibits two extreme cases. On one hand, when latency is minimal, further reductions in delay do not significantly enhance QoS. On the other hand, excessive delays become intolerable, leading to a substantial degradation in QoS to unacceptable levels. In order to approximate this trend and jointly optimize it with the immersive experience quality below, we refer to [28] and convert it into a scoring model. The transmission delay score of user $u$ at time slot $t$ can be expressed as
$F_{u,t}^{T}=\begin{cases}\frac{1}{1+e^{-\zeta(T_{{\max}}-T_{u,t})}},&T_{u,t}\leq T_{\max}\\ 0,&T_{u,t}\geq T_{\max}\end{cases},$ (27)
where $T_{{\max}}$ is the the maximized tolerable transmission delay. $\zeta$ is a constant deciding the steepness of the satisfactory curve.
The quality of the immersive experience is also one of the key factors affecting users’ QoS. In the considered scenario of panoramic video transmission, the immersive experience quality $Q_{u,t}$ is related to the semantically encoded CBR and the user’s SNR, i.e,
$Q_{u,t}=f_{\text{APVST}}\left(o_{u,t},\vartheta_{u,t}\right)$ (28)
where $f_{\text{APVST}}$ means the function of the APVST network. $Q_{u,t}$ is chosen as WS-PSNR or WS-SSIM. $o_{u,t}$ and $\vartheta_{u,t}$ are CBR and SNR for user $u$ at time slot $t$, respectively. To accelerate the optimization of the problem, we employ polynomial regression, nonlinear regression, and random forests to approximate the fitting of $f_{\text{APVST}}$.
Similar to transmission delay, there are two extremes in immersive experience quality. In order to perform joint optimization with the above transmission delay, it is converted into a score through maximum-minimum quantization. The immersive experience quality score $S_{u,t}^{Q}$ for user $u$ at time slot $t$ can be expressed as
$F_{u,t}^{Q}=\begin{cases}1,&Q_{u,t}\geq Q_{\max}\\ \frac{Q_{u,t}-Q_{\text{min}}}{Q_{\max}-Q_{\text{min}}},&Q_{\text{min}}<Q_{u,t}<Q_{\max}\\ 0,&Q_{u,t}\leq Q_{\text{min}}\end{cases},$ (29)
where $Q_{\text{min}}$ and $Q_{\max}$ are the upper and lower limits of immersive experience quality, respectively. Then the RSMA-enabled semantic information transmission optimization problem can be expressed as
$\max_{\mathbf{p}_{t},\mathbf{o}_{t},\mathbf{C}_{t}}\quad$ $\sum_{u=1}^{\left|\mathcal{U}\right|}F_{u,t}=\sum_{u=1}^{\left|\mathcal{U}\right|}\left[\kappa F_{u,t}^{T}+(2-\kappa)F_{u,t}^{Q}\right]$ (30)
s.t. $p_{t}^{c}+\sum_{u=1}^{\left|\mathcal{U}\right|}p_{u,t}^{p}\leq P_{{\max}},$ (30a)
$0<o_{u,t}\leq o_{\max},$ (30b)
$\sum_{u=1}^{\left|\mathcal{U}\right|}C_{u,t}=R_{t}^{c},$ (30c)
where $\mathbf{o}_{t}=\left[o_{1,t},\ldots,o_{u,t},\ldots,o_{\left|\mathcal{U}\right|,t}\right]$ is the CBR verctor at time slot $t$, and $\mathbf{C}_{t}=\left[C_{1,t},\ldots,C_{u,t},\ldots,C_{\left|\mathcal{U}\right|,t}\right]$ is the common rate allocation vector at time slot $t$. $\kappa$ is the balance coefficient between the transmission delay score and the immersive experience quality score. In the above optimization problem, constraint (30a) means that the power of the precoding matrix at time slot $t$ can’t exceed the maximum transmit power. Constraint (30b) means the CBR can’t exceed the maximum of CBR $o_{\max}$, where $o_{\max}$ is decided by quantized set $\mathcal{Q}$. (30c) is common rate allocation constraint to ensure that the common message stream can be successfully decoded by all users.
We can see that problem (30) is a mixed continuous-discrete non-convex optimization problem. Additionally, the time-varying nature of the channel and the variations in the user’s FoV further complicate the resolution of the problem. The solutions proposed by [10] are not applicable to our considered scenario. This is due to the fact that these approaches do not take into account the temporal impacts on the environment, nor do they optimize the transmission of information at a semantic level. Direct application of these solutions to our scenario may lead to local optima, thereby degrading the system’s transmission performance and affecting the users’ QoS. Considering the continuity of time, and since the transmitted information is video, there exists a certain semantic correlation between consecutive frames. Moreover, the user’s FoV continuously moves over time. Based on these analyses, we can transform the problem into a MDP. By making appropriate decisions at different time instances, we aim to maximize the total user score. Furthermore, we propose a solution based on PPO-based algorithm, the specifics of which will be detailed in Section IV.
(a) Network structure of semantic transmitter
(b) Network structure of semantic receiver
Figure 4: Network structures of the primary link in APVST. $k\times k$ Conv is a convolution with $k\times k$ filters, and the output channels of convolution are given on horizontal line. $\uparrow 2$ and $\downarrow 2$ indicate upsampling and downsampling with a stride of 2. GDN denotes the Generalised Divisive Normalization in [29], IGDN denotes the inverse operation of GDN.
III Internal Structure and Training Procedure of APSVT
@@
In this section, we introduce the internal structure and training procedure of the APVST network. The disassembled parts of the internal structure are shown in Fig. 4 and Fig. 5. Since the structures of primary and motion links are similar, the primary link will be introduced in detail.
III-A The Overall Structure of APSVT